### v1

# Gauge-Mediated Contagion: A Quantum Electrodynamics-Inspired Framework for Non-Local Epidemic Dynamics and Superdiffusion

Abstract
[…]
I Introduction
The mathematical modeling of infectious diseases has traditionally relied on the classical Kermack-McKendrick SIR equationsKermack and McKendrick (1927); Anderson and May (1991); Murray (2002), which assume a well-mixed population and instantaneous local contact. These deterministic differential equations provide a very useful method to study epidemics. With the increase in epidemic data, non-linear behavior, such as long-range spatial correlation have been observed in real-world pandemic data. Other phenomena, such as “burst” dynamics, multi-focal outburst, or super-diffusion, cannot be explained only by local interactionBrockmann et al. (2006); Brockmann and Helbing (2013); Vazquez et al. (2007); González et al. (2008). So, to be able to explain them, non-local interactions has to be assumed. Different methods have been proposed to introduce non-local characteristics in $SIR$-like model as, for instance, fractional derivatives or Levy-flight modelsAngstmann et al. (2017); Janssen et al. (1999).
In this study, we propose a unified framework inspired by Quantum Electrodynamics (QED). We treat the pathogen as a mediator field $\varphi(x,t)$ that exists independently of the hosts and carries the “epidemic charge” between susceptible ($\phi_{S}$) and infected ($\phi_{I}$) matter fields. By applying the Doi-Peliti second-quantization formalism, we map $SIR$-like models onto a continuous field theory. In this formalism, the dynamics are given by a generalized mean action principle, and in the “classical” limit, the $SIR$ equations are obtained. The QED techniques can be used to compute any observables through the use of Feynman rules applied to our action. The non-local characteristics naturally appear in this model once integrating out the pathogen fields. We show that the Vacuum polarization, which describes in our model how a susceptible population responds and “dressed” the pathogen field mass, modifying its mass $m_{R}$ (which is related to its propagation length ($\xi\approx 1/m_{R}$)). We derive the Effective Reproductive Number ($R_{eff}$) as a renormalized coupling constant and show how Debye Screening fundamentally alters the perceived infectivity across different scalesMollison (1977); Pastor-Satorras et al. (2015); Brockmann and Helbing (2013). The well-known epidemic threshold ($R_{0}=1$) is re-formulated as a symmetry-breaking phase transition where the effective pathogen mass vanishes ($m_{R}^{2}\to 0$)Grassberger (1983); Cardy and Grassberger (1985); Täuber (2014); Janssen et al. (2004); Hinrichsen (2000). This situation is very well-know in QED as when $m^{2}_{eff}\leq 0$, it describes a symmetry breaking phase transitionWeinberg (1996); Landau and Lifshitz (2013); Zinn-Justin (2021); Peskin and Schroeder (1995). Furthermore, we analyze the impact of super-spreading hosts, demonstrating how they can be easily introduced in this model, and they will affect the $R_{eff}$, for instance. Finally, we propose a comprehensive operational workflow to illustrate how our model can be calibrated with real-world clinical and environmental data to inform policy decisions.
The paper is organized as follows: Section II defines the field definitions and the action of the QED-inspired $SIR$ model. Section III derives the generating functional and the effective interaction kernel. Section IV details the computation of vacuum polarization and mass renormalization. Section V discusses the implications of Debye screening on the epidemic threshold. Section VI computes the renormalized $R_{\text{eff}}$ via vertex corrections. Section VII examines the impact of super-spreaders on the critical point. Section VIII summarizes the main results. In Section IX, we present the workflow for practical implementation. We conclude in Section X by suggesting future extensions into Non-Abelian gauge theories to model viral strain competition.
II Theoretical framework
II.1 The Action and Field Definitions
We define two complex ”matter” fields in the Doi-Peliti sensePeliti (1985); Doi (1976); Täuber (2014), representing Susceptibles ($\phi_{S},\hat{\phi}_{S}$) and Infecteds ($\phi_{I},\hat{\phi}_{I}$), and an auxiliary scalar field $\varphi(x,t)$ representing the pathogen concentration in the environment (the ”gauge” mediator). The total action $S$ is the sum of the population dynamics and the mediator field:
$S=\int d^{d}xdt\left[\sum_{a=S,I}\hat{\phi}_{a}\partial_{t}\phi_{a}+\frac{1}{2}\varphi\left(\partial_{t}-D_{\varphi}\nabla^{2}+m_{0}^{2}\right)\varphi+\mathcal{H}_{int}\right]$ (1)
where $\mathcal{D}^{-1}=(\partial_{t}-D_{\varphi}\nabla^{2}+m_{0}^{2})$ is the inverse propagator of the pathogen with bare mass $m_{0}$ (decay rate) and
  • $\phi_{a}(x,t)$ (Density Fields): These are the bosonic coherent state fields for the compartments $a\in\{S,I\}$. $\phi_{S}$ represents the density of susceptibles and $\phi_{I}$ the density of infecteds at spacetime point $(x,t)$.
  • $\hat{\phi}_{a}(x,t)$ (Response Fields): Also known as the auxiliary or ”conjugate” fields in the Doi-Peliti formalism. They track the stochastic fluctuations and the conservation of probability. $\hat{\phi}=1$ corresponds to the deterministic (mean-field) limit.
  • $\varphi(x,t)$ (Pathogen Field): The mediator field. It represents the local concentration of the pathogen (e.g., viral load in the air or water) that exists independently of the hosts.
  • $\mathcal{H}_{int}$ (Interaction Hamiltonian): Describes the coupling between the hosts and the pathogen field. In this model, it typically takes the form of Yukawa-type coupling: $g\phi\varphi$
  • The term $D_{\varphi}\nabla^{2}$ represents the Laplacian operator scaled by the diffusion coefficient. In the path integral, this term acts as the ”spatial penalty”: it ensures that spatial gradients in the pathogen field are smoothed out over time.
  • $d$ represents the spatial dimensionality of the system.
The inverse propagator is derived from the equation of motion of the pathogen in a vacuum (i.e., without host interference). We assume the pathogen undergoes standard diffusion and exponential decay. If we assume the pathogen reaches a steady state faster than the host population changes ($\partial_{t}\approx 0$), the spatial propagator becomes:
$\mathcal{D}(r)\propto\frac{1}{D_{\varphi}}\frac{e^{-r\sqrt{m_{0}^{2}/D_{\varphi}}}}{r^{d-2}}$ (2)
For $d=3$, the Yukawa potential is recuperated. To derive the SIR equations within the QED-inspired field theory, the interaction Hamiltonian $\mathcal{H}_{int}$ must be constructed using the Doi-Peliti operators. The interaction Hamiltonian is divided into the Contagion term and the Recovery term:
$\mathcal{H}_{int}=\mathcal{H}_{contagion}+\mathcal{H}_{recovery}$ (3)
[…]
$\mathcal{H}_{contagion}=\beta\int d^{d}x\left(\hat{\phi}_{S}\hat{\phi}_{I}-\hat{\phi}_{I}^{2}\right)\phi_{S}\varphi$ (4)
  • $\phi_{S}\varphi$: The rate depends on the density of susceptibles and the local concentration of the pathogen.
  • $(\hat{\phi}_{S}-\hat{\phi}_{I})\hat{\phi}_{I}$: This operator structure ensures that for every susceptible removed (via $\hat{\phi}_{S}$), an infected is added (via $\hat{\phi}_{I}$). The $\hat{\phi}_{I}^{2}$ term accounts for the ”birth” of a new infected in the presence of the existing infectious load.
The Recovery Term describes the transition from Infected to Removed ($I\to R$):
$\mathcal{H}_{recovery}=\gamma\int d^{d}x(\hat{\phi}_{I}-1)\phi_{I}$ (5)
where $(\hat{\phi}_{I}-1)$ is this operator removes an infected individual from the ”active” field. The 1-1 term is required for the normalization of the probability distribution in the Fock space.
To fully close the SIR loop, we must also define how the pathogen field $\varphi$ is produced. This is usually added as a linear coupling in the action:
$\mathcal{L}_{production}=g\hat{\varphi}\phi_{I}$
This term ensures that the infected population $I$ acts as a current source for the pathogen field $\varphi$, analogous to how a moving charge produces an electromagnetic field in QED.
To recover the $SIR$ model as the “classical” limit, one has to obtain the saddle point of the Action.
The path integral is defined as $\mathcal{Z}=\int\mathcal{D}\phi\mathcal{D}\hat{\phi}\,e^{-S[\phi,\hat{\phi}]}$. The saddle point is found by taking the functional derivatives:
$\frac{\delta S}{\delta\hat{\phi}_{a}}=0\quad\text{and}\quad\frac{\delta S}{\delta\phi_{a}}=0$
where $a=S$ and $I$. As proposed in Doi-Peliti formalism,when $\hat{\phi}=1$, the entire interaction Hamiltonian $\mathcal{H}_{int}$ becomes zero. As a result, The equations of motion for the “real” fields $\phi_{a}$ (the densities) reduce to the deterministic rate equations (the SIR ODEs). In this QED-inspired model and within the Doi-Peliti formalism:
  • $\phi_{a}$ describes the mean density of individuals.
  • $(\hat{\phi}_{a}-1)$ describes the fluctuations (stochastic noise) away from that mean.
So, to get $SIR$ equations with this $\mathcal{H}_{int}$, we evaluate the Hamilton-Jacobi equations at the physical saddle point where the response fields $\hat{\phi}_{S}=1$ and $\hat{\phi}_{I}=1$:
$\dot{S}=\frac{\partial\mathcal{H}_{int}}{\partial\hat{\phi}_{S}}=\beta\phi_{S}\varphi\hat{\phi}_{I}\to-\beta S\varphi$
$\dot{I}=\frac{\partial\mathcal{H}_{int}}{\partial\hat{\phi}_{I}}=\beta\phi_{S}\varphi(2\hat{\phi}_{I}-\hat{\phi}_{S})-\gamma\phi_{I}\to\beta S\varphi-\gamma I$
The resulting “QED-SIR” system of equations is then given as
$\frac{\partial S}{\partial t}$ $=$ $-\beta S\varphi$ (6)
$\frac{\partial I}{\partial t}$ $=$ $\beta S\varphi-\gamma I$ (7)
$(\frac{\partial}{\partial t}-D_{\varphi}\nabla^{2}+m_{0}^{2})\varphi$ $=$ $gI$ (8)
The standard SIR model is a ”point-like” approximation of this theory. To recover it, we assume the pathogen field reaches equilibrium instantaneously and has no spatial diffusion ($D_{\varphi}\to 0$), one gets $m_{0}^{2}\varphi=gI\implies\varphi=\frac{g}{m_{0}^{2}}I$.
$\frac{dS}{dt}=-\left(\frac{\beta g}{m_{0}^{2}}\right)SI$
$\frac{dI}{dt}=\left(\frac{\beta g}{m_{0}^{2}}\right)SI-\gamma I$
By defining the effective transmission rate as $\beta_{eff}=\frac{\beta g}{m_{0}^{2}}$, we exactly recover the classical Kermack-McKendrick SIR equations.
II.2 Generating Functional of the gauge mediated SIR Model
Because the pathogen field $\varphi$ is Gaussian in the action, we can integrate it out to find the effective interaction between individuals. We use the Gaussian integral identity $\int e^{-\frac{1}{2}\varphi A\varphi+J\varphi}\propto e^{\frac{1}{2}JA^{-1}J}$ The source $J$ for the pathogen field is the infected population density $g_{1}\phi_{I}$. The effective interaction term becomes:
$S_{eff}=\int dx\ dy\ dt\left[(g_{1}\phi_{I}(x,t))D(x-y)(g_{2}\phi_{S}(y,t))\right]$ (9)
where D(xy)D(x-y) is the Green’s Function of the pathogen field:
$D(r)=\frac{e^{-m_{0}r}}{r^{d-2}}$ (10)
Integrating out the mediator field, a non-local action is generated. Without the field $\varphi$, the interaction term in the SIR model changes:
  • Before: Local interaction with a field: $gS(x)\varphi(x)$.
  • After: Non-local interaction between people: $\int dy\,S(x)\mathcal{K}(x-y)I(y)$
The Kernel $\mathcal{K}(x-y)$ is the Green’s function (propagator) of the pathogen. This proves that ”non-locality” in epidemics isn’t just a mathematical trick; it is a direct consequence of a mediator (the pathogen) moving through space. If the propagator of the field you integrated out has specific scaling properties (for example, if the pathogen moves via a ”turbulent” or ”fractional” diffusion), the resulting kernel $\mathcal{K}$ will naturally be a Lévy distribution.
Table 1: Mapping between Quantum Electrodynamics (QED) and the Proposed Epidemic Gauge Theory (QED-SIR).
Feature Quantum Electrodynamics (QED) Epidemic Gauge Theory (QED-SIR)
Matter Field Electron / Fermion ($\psi$) Host Population ($\phi_{S},\phi_{I}$)
Gauge Field Photon ($A_{\mu}$) Pathogen Field ($\varphi$)
Interaction Minimal Coupling (Vertex) Gauge-mediated Contagion
Coupling Constant Electric Charge ($e$) Epidemic Charge ($g,\beta$)
Propagator Massless ($1/k^{2}$) Massive ($1/(k^{2}+m_{0}^{2})$)
Interaction Range Infinite (Coulombic) Finite (Yukawa / Screened)
Fundamental Symmetry $U(1)$ Local Phase Invariance Probability Conservation / Shift Invariance
The statistical properties of the epidemic are encapsulated in the generating functional $\mathcal{Z}[J,\eta]$, defined asAltland and Simons (2010); Weinberg (1995); Zinn-Justin (2021); Peskin and Schroeder (1995):
$\mathcal{Z}[J,\eta]=\int\mathcal{D}[\Phi]\exp\left(-S[\Phi]+\int d^{d}xdt\mathcal{J}\cdot\Phi\right)$ (11)
By performing the Gaussian integration over the mediator field $\varphi$, we obtain the effective generating functional for the host densities:
$\mathcal{Z}_{eff}[J]=\int\mathcal{D}[\phi,\hat{\phi}]\exp\left(-S_{0}+\frac{1}{2}\int\rho_{eff}\mathcal{D}(x-y)\rho_{eff}dy\right)$ (12)
where $\rho_{eff}=g\phi_{I}-\beta\phi_{S}(\hat{\phi}_{I}-\hat{\phi}_{S})\hat{\phi}_{I}$. For a non-linear model like SIR, we cannot solve the remaining integral over $\phi$ exactly. We expand the generating functional around the Gaussian fixed point (the free theory) using a series of Feynman diagrams:
$\mathcal{Z}_{eff}[J]\approx\exp\left(\int\mathcal{L}_{int}\left(\frac{\delta}{\delta J}\right)\right)\mathcal{Z}_{0}[J]$ (13)
Where $\mathcal{Z}_{0}$ is the ”free” generating functional (only recovery and diffusion, no contagion).
II.3 Feynman Rules and Diagrammatic Representation
The exponential operator in Eq.(13) can be expanded in a Taylor series, generating the perturbative expansion of the theory. Each term in this series corresponds to a specific number of interaction events (vertices) occurring in spacetime:
$\mathcal{Z}[J]=\left[1-\int d^{d}x\mathcal{L}_{int}\left(\frac{\delta}{\delta J}\right)+\frac{1}{2!}\left(\int d^{d}x\mathcal{L}_{int}\left(\frac{\delta}{\delta J}\right)\right)^{2}-\dots\right]\mathcal{Z}_{0}[J]$ (14)
The evaluation of these terms relies on Wick’s Theorem. The functional derivatives contained in $\mathcal{L}_{int}$ act on the Gaussian functional $\mathcal{Z}_{0}[J]$, effectively pairing the field operators. Mathematically, each pair of functional derivatives brings down a free propagator $\Delta_{F}(x-y)$, which represents the inverse of the differential operator in the free action $S_{0}$.
This mathematical structure admits a direct graphical representation known as Feynman diagrams:
  • Each integration variable $\int d^{d}x$ corresponds to a vertex in the diagram, representing a local interaction event (e.g., a susceptible host absorbing a pathogen).
  • Each contraction $\frac{\delta}{\delta J(x)}\frac{\delta}{\delta J(y)}$ corresponds to a line (propagator) connecting points $x$ and $y$, representing the causal propagation of a field (e.g., the pathogen moving through the environment).
  • The pre-factors (like $1/n!$) correspond to the symmetry factors of the diagram.
The physical observables can be easily computed expanding the generating functional using Feynman diagrams techniques:
  1. 1.
    The Mean Density (1-point function).The classical SIR trajectory is found by the first derivative:
    $\langle\phi_{I}(x,t)\rangle=\left.\frac{\delta\ln\mathcal{Z}}{\delta J_{I}(x,t)}\right|_{J=0}$
  2. 2.
    The Epidemic Correlation (2-point function): to see how city A affects city B, we take the second derivative:
    $\langle\phi_{I}(x_{1})\phi_{I}(x_{2})\rangle=\left.\frac{\delta^{2}\ln\mathcal{Z}}{\delta J_{I}(x_{1})\delta J_{I}(x_{2})}\right|_{J=0}$
Table 2: Feynman Rules for the QED-inspired SIRSIR Stochastic Field Theory.
Component Mathematical Expression Diagrammatic Symbol
Infected Propagator $G_{I}(k,\omega)=(-i\omega+D_{I}k^{2}+\gamma)^{-1}$ Solid Line
Pathogen Propagator $D_{\varphi}(k,\omega)=(-i\omega+D_{\varphi}k^{2}+m_{0}^{2})^{-1}$ Wavy Line
Emission Vertex $g$ Solid-Wavy Junction
Infection Vertex $\beta$ Solid-Dashed-Wavy Node
Loop Integral $\int\frac{d^{d}q}{(2\pi)^{d}}\frac{d\Omega}{2\pi}$ Closed Loop
III Vacuum Polarization (the self-energy $\Pi$ of the pathogen field)
Vacuum Polarization represents the way a population “responds” to the presence of a pathogen, effectively altering the transmission environment.When a pathogen moves through a vacuum of susceptibles, it doesn’t just sit there; it constantly couples to the hosts. An infected person shedding a virus (the mediator boson) creates a local “cloud” of potential new infections.This interaction modifies the propagation of the field modifying its mass. So, one expect that instead of just decaying according to its biological half-life ($m_{0}$), the pathogen’s reach is modified by the density of available hosts.
[…]
Figure 1: Vacuum Polarization of the Pathogen Field. The one-loop self-energy diagram $\Pi(q)$ for the gauge mediator $\varphi$. The incoming pathogen (wavy line) fluctuates into a virtual pair of Susceptible ($\phi_{S}$) and Infected ($\phi_{I}$) matter fields before recombining. This process is responsible for the renormalization of the pathogen mass $m_{R}$ and the emergence of the Debye screening length $\lambda_{D}$ in the effective potential.
The self-energy is the mathematical value of the loop. According to the Feynman rules, we integrate over the internal momentum $q$ and frequency $\Omega$:
$\Pi(k,\omega)=g\beta\int\frac{d^{d}q}{(2\pi)^{d}}\int\frac{d\Omega}{2\pi}G_{I}(q,\Omega)G_{S}(k-q,\omega-\Omega)$ (15)
where assuming a static background of susceptibles $S_{0}$ (the ”vacuum” density):
  • $G_{S}\approx S_{0}/(-i\Omega+\epsilon)$
  • $G_{I}=1/(-i(\omega-\Omega)+D_{I}(k-q)^{2}+\gamma)$
Using the residue theorem for the $\Omega$ integral, we pick up the pole from the susceptible line. In the static limit ($\omega\to 0,k\to 0$):
$\Pi(0,0)\approx\frac{g\beta S_{0}}{\gamma}$
The full propagator $D_{dressed}$ is found by summing the geometric series of these bubbles (the Dyson Equation):
$D_{dressed}(k)=\frac{1}{D_{\varphi}k^{2}+m_{0}^{2}-\Pi(0,0)}$
The denominator now defines the Renormalized Mass $m_{R}$:
$m_{R}^{2}=m_{0}^{2}-\frac{g\beta S_{0}}{\gamma}$ (16)
If the population density $S_{0}$ is high enough such that $\frac{g\beta S_{0}}{\gamma}>m_{0}^{2}$, the mass squared becomes negative.A negative $m_{R}^{2}$ in field theory signals a symmetry breaking or vacuum decay. In epidemiology, this is exactly the point where $R_{0}>1$, and the disease-free equilibrium becomes unstable, leading to a spontaneous pandemic. The interaction with the population increases the effective range of the virus ($\xi=1/m_{R}$). The ”vacuum” helps the pathogen travel further than it could in a sterile environment.
IV $R_{0}$ in QED-SIR model
The Reproductive Number $R_{0}$ is defined as the ratio of the ”production/infection” rate to the ”decay/removal” rate. In our gauge model, this is the ratio of the Population Polarization to the Bare Environmental Mass:
$R_{0}=\frac{\text{Field Coupling Strength}}{\text{Environmental Resistance}}=\frac{g\beta S_{0}}{\gamma m_{0}^{2}}$ (17)
where:
  • $g$ (Emission): Rate at which an infected person ”charges” the environment with virus.
  • $\beta$ (Absorption): Cross-section of a susceptible person catching the virus from the field.
  • $S_{0}$ (Density): The ”Dielectric Constant” of the vacuum (the available susceptible host density).
  • $\gamma$ (Host Recovery): The rate at which the ”Matter Field” sources are removed.
  • $m_{0}^{2}$ (Pathogen Decay): The ”Mass” that limits the range of the gauge boson.
V Debye screening and $R_{0}$
In our QED-inspired $SIR$ model, Debye Screening is the process by which the susceptible population absorbs and shields the pathogen field, preventing it from reaching distant hosts.
In a sterile environment, a pathogen might travel far. In a crowded population, the matter fields (people) interact with the mediator field (pathogen), effectively increasing its mass and shortening its range. We start with the static classical equation for the pathogen field $\varphi$ in the presence of a host population. From our previously derived equations of motion:
$(\nabla^{2}-m_{0}^{2})\varphi=-gI$ (18)
Where $m_{0}$ is the bare environmental decay and $gI$ is the source from infected individuals. In a susceptible population $S_{0}$, the pathogen induces a ”polarization.” If the field $\varphi$ is present, it creates a change in the infection rate. Near the threshold, we can linearize the response of the ”Infected” density $I$ to the field $\varphi$:
$\delta I\approx\frac{\beta S_{0}}{\gamma}\varphi$ (19)
This represents the susceptible individuals becoming polarized into a pre-infectious state by the field.
Substitute the response back into the field equation :
$\nabla^{2}\varphi-m_{0}^{2}\varphi=-g\left(\frac{\beta S_{0}}{\gamma}\varphi\right)$
Rearrange the terms to group the $\varphi$ coefficients:
$\nabla^{2}\varphi-\left(m_{0}^{2}-\frac{g\beta S_{0}}{\gamma}\right)\varphi=0$
The coefficient of $\varphi$ defines the Inverse Screening Length Squared ($\kappa^{2}$). We define the Renormalized Mass $m_{R}^{2}$ as:
$m_{R}^{2}=m_{0}^{2}\left(1-\frac{g\beta S_{0}}{\gamma m_{0}^{2}}\right)$
Using our definition of $R_{0}=\frac{g\beta S_{0}}{\gamma m_{0}^{2}}$, we get:
$m_{R}^{2}=m_{0}^{2}(1-R_{0})$
The Debye Screening Length is the characteristic distance the pathogen can travel before being ”screened” out by host interactions:
$\lambda_{D}=\frac{1}{m_{R}}=\frac{1}{m_{0}\sqrt{1-R_{0}}}$
Debye screening does not just change the distance of infection; it fundamentally alters the effective $R_{0}$ perceived at different scales. Because of the screening, the interaction potential $V(r)$ shifts from a Yukawa potential ($V(r)\sim\frac{e^{-r/\lambda_{D}}}{r^{d-2}}$ to a long range propagator.
As $R_{0}\to 1$ from below:
$\lambda_{D}\to\infty$
This is the Critical Opalescence of an epidemic. The screening fails, the pathogen becomes ”massless” (long-ranged), and $R_{0}$ effectively becomes global.
VI $R_{eff}$ Computation
In our QED-inspired $SIR$ model, the Effective Reproductive Number $R_{eff}$ is the renormalized version of $R_{0}$.The first correction to $R_{0}$ is the reduction of the available “charge” (susceptibles) as the epidemic progresses. In the saddle-point approximation ($\hat{\phi}=1$), one obtains the classical result where $R_{eff}$ drops.:
$R_{eff}(t)=\frac{g\beta S(t)}{\gamma m_{0}^{2}}=R_{0}\frac{S(t)}{S_{0}}$ (20)
As we derived in the previous section, the pathogen doesn’t interact with a single person; it interacts with a screened cloud. The effective transmission is modified by the Renormalized Propagator at zero momentum:
$R_{eff}^{(screened)}=R_{0}\times\frac{m_{0}^{2}}{m_{R}^{2}}$ (21)
Substituting our result $m_{R}^{2}=m_{0}^{2}(1-R_{0}\frac{S}{S_{0}})$:
$R_{eff}\approx\frac{R_{0}\frac{S}{S_{0}}}{1-\text{Correction Term}}$
This reflects how the “reach” of each infected individual is boosted by the proximity of other susceptible hosts.
Using the Feynman rules, the 1-loop correction to the infection vertex involves a pathogen line and an infected line:
$\beta_{ren}=\beta_{0}\left(1-\frac{\beta_{0}}{\gamma}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{D_{\varphi}k^{2}+m_{0}^{2}}\right)$
Evaluating the integral in $d=2$:
$\beta_{ren}\approx\beta_{0}\left(1-\frac{\beta_{0}}{4\pi\gamma D_{\varphi}}\ln\left(\frac{\Lambda}{m_{R}}\right)\right)$
Where $\Lambda$ is the ultraviolet cutoff (the inverse of the minimum distance between people).
[…]
Figure 2: 1-Loop Vertex Correction. The Feynman diagram representing the renormalization of the infection rate $\beta$. The main interaction (center vertex) is ”dressed” by a virtual pathogen loop (wavy line) carrying momentum $k$. This loop integral generates the logarithmic correction $\ln(\Lambda/m_{R})$ in two dimensions, effectively screening the coupling constant at large distances.
Combining the depletion of hosts and the renormalization due to spatial fluctuations, the full $R_{eff}$ for your model is:
$R_{eff}(t)=R_{0}\frac{S(t)}{S_{0}}\left[1-\frac{\beta}{4\pi\gamma D_{\varphi}}\ln\left(\frac{\Lambda}{m_{R}(t)}\right)\right]$
The logarithmic term (the ”Fluctuation Correction”) is negative. This means that in a spatial world, $R_{eff}$ is always lower than the mean-field prediction. This is because infected individuals tend to be surrounded by “already-infected” or “recovered” people, shielding the susceptible vacuum. As the system approaches the peak, $m_{R}\to 0$, causing the logarithmic term to grow. This mathematically explains why the transition near the peak of a pandemic is ”slower” than simple SIR models suggest.
VII Superspreading hosts
We divide the infected population into two species: $I_{L}$ (low-emitters) and $I_{H}$ (high-emitters/super-spreaders). In QED, it will be equivalent to introduce charged particles with an electric charge bigger than electron. The pathogen field $\varphi$ is now sourced by a multi-component current:
$\mathcal{J}(x,t)=g_{L}I_{L}(x,t)+g_{H}I_{H}(x,t)$ (22)
To formalize it in our model, we define a population where a fraction $p_{i}$ has a charge $g_{i}$; the total vacuum polarization is the sum of all possible host-loop contributions. Each ”species” of host (low-emitters vs. super-spreaders) creates its own virtual ”bubble” in the propagator. The total self-energy $\Pi_{total}$ is:
$\Pi_{total}(k,\omega)=\sum_{i}p_{i}\Pi_{i}(k,\omega)$ (23)
The integral for the self-energy at zero momentum ($k\to 0,\omega\to 0$) becomes:
$\Pi(0,0)=\int P(g)dg\left[\frac{g\beta S_{0}}{\gamma}\right]$ (24)
where $P(g)$ describes this statistical distribution across the host population of people with a charge $g$.If $P(g)$ is a Delta function ($\delta(g-g_{0})$), everyone is equally infectious (Standard SIR model). If $P(g)$ is a Gamma or Power-law distribution (fat-tailed), it means a small number of individuals have a very large $g$ (Super-spreaders). This would make the integral dominated by the tail of the distribution, significantly increasing the effective coupling and the risk of vacuum instability ($m_{R}^{2}\to 0$).
The dressed pathogen mass $m_{R}$ is corrected by the average polarization of the vacuum:
$m_{R}^{2}=m_{0}^{2}-\frac{\beta S_{0}}{\gamma}\langle g\rangle$ (25)
where $\langle g\rangle=\int gP(g)dg$ is the First Moment of the Distribution, also called the expectation value of $g$. Eq.(25) is valid at tree-level. Including the one loop vaccum corrections as done in previous sections, one gets a correction that behaves as $\langle g^{2}\rangle$. Even if the average shedding $\langle g\rangle$ is low, a “fat tail” in $P(g)$ (large variance) will make $\langle g^{2}\rangle$ very large. This permits us to introduce the definition of the Heterogeneity Factor $\kappa$:
$\kappa=\frac{\langle g^{2}\rangle}{\langle g\rangle^{2}}$ (26)
which means that if $\kappa=1$, we have a homogeneous population (Standard SIR) ), whereas if $\kappa\gg 1$, we have a superspreading population. The Debye screening length becomes sensitive to the higher moments of the pathogen emission distribution. The effective Debye length $\lambda_{D}^{eff}$ expands as follows:
$\lambda_{D}^{eff}=\frac{1}{m_{0}\sqrt{1-R_{0}(1+\kappa)}}$ (27)
Their effects are the following:
  • The presence of super-spreaders (high $\kappa$) causes the correlation length to diverge ($\lambda_{D}^{eff}\to\infty$) much earlier than predicted by the standard, homogeneous $SIR$ model.
  • This mathematically explains why “giant clusters” or multi-focal outbreaks appear even when the population average $R_{0}$ suggests the epidemic is under control.
VIII Results and discussion
By treating the pathogen as a mediator boson, we can explain several “long-tail” and non-linear behaviors that classical SIR models fail to capture:
  1. 1.
    Ability to detect the onset of a pandemic: Our model naturally incorporates the concept of critical opalescence, a phenomenon from the physics of second-order phase transitions. This provides an early warning signal for impending epidemic waves. In the stable regime ($R_{eff}<1$), the pathogen field $\varphi$ is massive, and infection fluctuations are suppressed by the exponential decay of the propagator (Yukawa regime). However, as the system approaches the critical threshold ($R_{eff}\to 1$), the renormalized mass $m_{R}$ vanishes, leading to a divergence in the correlation length $\xi\propto 1/m_{R}$. In epidemiological data, this manifests as critical opalescenceJanssen et al. (2004); Cardy and Grassberger (1985): a sudden, dramatic increase in the variance of case counts and the emergence of self-similar infection clusters across multiple spatial scales, just before the onset of a pandemic. Through our mechanism of monitoring the Debye screening mass, this process can be studied in real time.
  2. 2.
    Super-spreaders as High-Intensity Gauge Sources: By introducing heterogeneity into the epidemic charge ($g$), we found that $R_{eff}$ is driven by the variance of the population rather than the mean. A small fraction of high-charge individuals (super-spreaders) can keep the epidemic vacuum unstable ($m_{R}^{2}<0$) even when the rest of the population is below the threshold. This explains why real-world pandemics often exhibit burst dynamics.
IX Operational Framework and Practical Application
Until now, we have shown how our QED-inspired model provides a very complete dynamical model of the disease transmission. So, the next step is to demostrate how in practical epidemic study, one can use it. Therefore in this section, we propose a bridge via a workflow diagram to connect the field theoretical formalism and public health policy. Unlike classical compartmental models that rely on static parameters, this model treats the environment as a dynamic mediator field, requiring a specific calibration of physical operators from clinical and environmental data that has to be monitored in real time.
IX.1 Calibration of Field Operators
The transition from a theoretical Lagrangian to a data-driven analysis requires mapping observable epidemiological quantities to the parameter $m_{0}$, $g$, and $\beta$ than appear in the action.
  1. 1.
    The Bare Mass ($m_{0}$) and Environmental Decay: The parameter $m_{0}$ defines the range of the pathogen in a vacuum (absence of hosts). It is physically determined by the environmental half-life ($t_{1/2}$) of the pathogen in aerosols or on surfaces for instance:
    $m_{0}\approx\sqrt{\frac{\ln 2}{D_{\varphi}t_{1/2}}}$ (28)
    where $D_{\varphi}$ is the diffusion coefficient of the viral vector (e.g., aerosolized droplets). Interventions such as UV-C lighting or HEPA filtration effectively increase $m_{0}$, shortening the interaction range.
  2. 2.
    Epidemic Charge ($g$) and Shedding Rates: The coupling constant $g$ represents the rate at which an infected host sources the pathogen field. This corresponds to the viral shedding rate (viral load) measured in clinical samples (e.g., RNA copies/mL).
  3. 3.
    Heterogeneity ($\kappa$) and Super-spreading: Recognizing that shedding rates are highly skewed, we define the heterogeneity parameter $\kappa$ based on the variance of the shedding distribution:
    $\kappa=\frac{\langle g^{2}\rangle}{\langle g\rangle^{2}}$ (29)
    A high $\kappa$ indicates the presence of super-spreaders.
IX.2 Operational Workflow
The application of the model follows a five-step workflow, illustrated in Figure 3.
  • Phase 1 (Parameterization): Clinical and environmental data are converted into the action parameters as explained above.
  • Phase 2 (Spatial Risk Mapping): The Debye screening length $\lambda_{D}$ is computed spatially. Regions where $\lambda_{D}$ diverges are identified and critical opalescence should be check to generate health public actions.
  • Phase 3 (Dynamic Tracking): The effective reproductive number $R_{eff}$ is monitored with 1-loop fluctuation corrections to account for spatial shielding effects.
  • Phase 4 (Stability Analysis): The system is monitored for ”Vacuum Instability” ($m_{R}^{2}\to 0$), a signal of critical opalescence preceding a phase transition.
  • Phase 5 (Mediator Shielding): Interventions are designed to modify the action parameters.
Clinical & Env. Data Calculate Parameters ($m_{0},g,\kappa$) Compute Debye Length $\lambda_{D}$ $\lambda_{D}\to\infty$? Local Clusters (Screened) Global Spread (Massless) Check Vacuum Stability ($m_{R}^{2}$) $m_{R}^{2}\approx 0$? Critical Opalescence Intervention: Mediator Shielding NoYesYesNo
Figure 3: Workflow for the practical application of the QED-SIR model. The process translates biological data into field-theoretic variables to assess the stability of the epidemiological vacuum.
X Conclusion and Future Perspectives
In this work, we have proposed a QED-inspired $SIR$ model. In this model, all the dynamics of the disease transmission is implicitly defined. By replacing the ”direct contact” paradigm of classical SIR models with a dynamic interaction mediated by a mediator field $\varphi$, we have demonstrated that QED mathematical techniques can be extended to this epidemiological model giving us mechanisms to describe various observables than the standard $SIR$ model fails to explain as non-local interactions, non-linear effects,the onset of a pandemic and variation of $R_{eff}$. Crucially, we show that the non-local interactions observed in real-world data are not ad-hoc additions without justifications more than to fit the data but emerge naturally from the theory by integrating out the pathogen field. We also have shown how naturally spatial correlations are emerging.
We have identified the epidemic threshold not only as a statistical parameter ($R_{0}$), but as a symmetry-breaking phase transition where the vacuum becomes unstable. This formalism offers a theoretical basis for “Pandemic Opalescence”, explaining the divergence of correlation lengths and the critical fluctuations observed prior to major outbreaks.
Furthermore, the application of QED techniques, specifically the use of Feynman diagrams,give us a powerful perturbative toolkit for calculating epidemiological observables. We have shown that host heterogeneity, particularly super-spreading events, can be naturally incorporated into this formalism as high-intensity gauge sources that renormalize the effective coupling.Finally, while the current model utilizes a $U(1)$-like Abelian symmetry to describe a single pathogen, real-world scenarios often involve multiple competing viral strains. A natural and promising extension of this work lies in the development of a Non-Abelian Gauge Theory (Yang-Mills SIR) to describe the complex, non-linear competition between variants and the dynamics of cross-immunity.